# Implications of recent Higgs data for the light Higgs bosons in the GNMSSM

### v2

@@ Abstract
This study systematically investigates the parameter space of the GNMSSM featuring light Higgs bosons, with a focus on the exotic decay of the 125 GeV Higgs boson ($h$) into pairs of the lightest CP-odd ($a_{1}$) or CP-even ($h_{1}$) Higgs bosons under constraints from recent Higgs data. We perform an initial scan of the parameter space under basic constraints from HiggsSignals-2.6.2 and HiggsBounds-5.10.2, followed by the application of comprehensive constraints using the HiggsTools framework. Comparative analysis reveals that HiggsTools imposes significantly more stringent constraints than its legacy components alone. Furthermore, its exclusion capability substantially surpasses that of ATLAS direct searches targeting individual decay channel.
By analyzing the distinct constraining effects of the subpackages HiggsSignals and HiggsBounds within HiggsTools, we find HiggsBounds constraints to be dominant across both the $h_{1}$ and $h_{2}$ scenarios. Nearly all parameter points excluded by HiggsSignals or direct LHC decay channel analyses are also excluded by HiggsBounds, attributable to the direct constraints HiggsBounds provides for non-SM Higgs searches, in contrast to the indirect constraints from HiggsSignals. In the $h_{2}$ scenario, HiggsSignals can exclude points with suppressed $Br(h\to a_{1}a_{1}\to\tau\tau bb)$ due to the kinematically accessible decay $h_{2}\to h_{1}h_{1}$, which exhibits large branching fractions and modifies SM-like Higgs properties. Our analysis shows that experimental data require the observed $h$ boson to be predominantly SM-like in both scenarios. Specifically, in the $h_{2}$ scenario, the lighter singlet-dominated state $h_{s}$ must possess high singlet purity and negligible SM-like components.
@@ 1 Introduction
The discovery of the 125 GeV Higgs boson at the Large Hadron Collider (LHC) [1, 2] in 2012 marked a monumental milestone in particle physics, confirming the mechanism of electroweak symmetry breaking (EWSB) as described by the Standard Model (SM). However, despite its tremendous success, the SM leaves several fundamental questions unresolved, such as neutrino masses, the hierarchy problem, and the dark matter problem. These unresolved issues have motivated extensive exploration of physics beyond the Standard Model (BSM), where extended Higgs sectors and new symmetries often predict additional scalar states.
@@
Supersymmetry (SUSY) stands out as a compelling BSM candidate, offering solutions like addressing the hierarchy problem and enabling gauge coupling unification. The Minimum Supersymmetric Standard Model (MSSM) [3, 4, 5, 6], as the most economical SUSY realization, has been extensively studied. Yet, it faces significant limitations from dark matter constraints and struggles with issues like the $\mu$ problem and the little hierarchy problem. As a natural and minimal extension, the Next-to-Minimal Supersymmetric Standard Model (NMSSM) [7, 8, 9] introduces a gauge-singlet Higgs superfield. Its scalar component acquires a non-zero vacuum expectation value (vev), dynamically generating an effective $\mu$-term. Crucially, mixing between the singlet and doublet superfields produces a new CP-even Higgs ($h_{s}$) that mixes with MSSM-like states. This mixing can suppress the lightest Higgs mass while allowing the next-to-lightest Higgs to match the observed 125 GeV signal [10, 11, 12, 13, 14, 15, 16]. This scenario alleviates fine-tuning and offers rich phenomenology, including Higgs-to-Higgs decays [17, 18, 19, 20, 21, 22, 23, 24] or exotic decays involving neutralinos [25].
However, the predictive power of the $Z_{3}$-invariant NMSSM is constrained by its superpotential structure, motivating generalizations like the General NMSSM (GNMSSM) [26, 27, 28, 29, 30, 31]. By adding extra singlet couplings or $Z_{3}$-symmetry-breaking terms, the GNMSSM enhances flexibility in the Higgs mass matrix, enabling lighter CP-even and CP-odd states consistent with LHC data. These light Higgs bosons could evade direct detection due to suppressed SM couplings but might appear through exotic decays of the 125 GeV Higgs $h$, such as $h\to h_{s}h_{s}/a_{1}a_{1}$ (where $a_{1}$ is the lightest CP-odd Higgs). Such rare decays transform the 125 GeV Higgs into a unique portal for discovering light hidden sector particles.
Both the ATLAS and CMS collaborations at the LHC have pursued extensive searches for light Higgs bosons through exotic decays of the 125 GeV Higgs boson [32], targeting final states like $4b$ [33], $\mu\mu bb$ [34, 35, 36], $\tau\tau bb$ [37, 38], $\gamma\gamma\tau\tau$ [39], and $4\mu$ [40, 41, 42, 43, 44]. Despite stringent limits from these searches, no significant excess has been observed, imposing strong constraints on the GNMSSM parameter space. This work primarily focuses on the promising yet challenging $\tau\tau bb$ final state [38]. Furthermore, we incorporate complementary constraints on the parameter space using the package HiggsTools [45], which combines and extends the functionalities of HiggsBounds and HiggsSignals. HiggsBounds tests BSM model compatibility against exclusion limits from new scalar searches, while HiggsSignals assesses predictions against experimental measurements of the 125 GeV Higgs boson production rates.
This study presents a comprehensive investigation of the 125 GeV Higgs boson’s exotic decays into pairs of lighter CP-even or CP-odd Higgs bosons within the GNMSSM framework. By systematically incorporating the stringent constraints from LHC direct Higgs searches and HiggsTools, we explore the viable parameter space capable of accommodating such signatures in the GNMSSM.
@@
This paper is structured as follows. Section 2 introduces the GNMSSM framework, detailing its Higgs sector and the key trilinear Higgs couplings essential for our analysis. In Section 3, we outline our scanning strategy and present the corresponding numerical results. Finally, Section 4 summarizes our conclusions.
@@ 2.1 Basics of the GNMSSM
The GNMSSM extends the MSSM by introducing a gauge-singlet superfield $\hat{S}$, carrying neither baryon nor lepton number. Consequently, the Higgs sector comprises not only the standard $SU(2)_{L}$ doublets $\hat{H}_{u}=(\hat{H}^{+}_{u},\hat{H}^{0}_{u})$ and $\hat{H}_{d}=(\hat{H}^{0}_{d},\hat{H}^{-}_{d})$, but also the singlet $\hat{S}$. Its gauge-invariant superpotential is given by  [8]:
@@
where $W_{\rm Yukawa}$ contains the MSSM quark and lepton Yukawa interactions. The dimensionless couplings $\lambda$ and $\kappa$ govern Higgs interactions, analogous to the $Z_{3}$-NMSSM. Crucially, the GNMSSM incorporates explicit $Z_{3}$-symmetry-breaking via the parameters $\mu$, $\mu^{\prime}$ and $\xi$. These terms resolve the tadpole problem [46, 8] and the cosmological domain-wall problem [47, 48, 49] inherent in the $Z_{3}$-invariant NMSSM. The parameter $\xi$ can be consistently eliminated by singlet field redefinition  [50], allowing us to set $\xi=0$ without loss of generality. Furthermore, $\mu$ and $\mu^{\prime}$ naturally arise from the spontaneous breaking of discrete $R$-symmetries ($\mathbb{Z}^{R}_{4}$ or $\mathbb{Z}^{R}_{8}$) at high scales [47, 51, 52, 50, 53]. These explicit $Z_{3}$-breaking terms significantly modify the neutral Higgs mass spectrum, generating phenomenology markedly richer than both the $Z_{3}$-NMSSM and the MSSM.
2.2 The Higgs Sector of GNMSSM
@@
with $H_{u},H_{d},$ and $S$ denoting the scalar components of the Higgs superfields, and $m^{2}_{H_{u}},m^{2}_{H_{d}},$ and $m^{2}_{S}$ being their supersymmetry-breaking masses. Following the electroweak symmetry breaking, the neutral components of the Higgs fields develop non-vanishing vevs:
@@
with $v=\sqrt{v_{u}^{2}+v_{d}^{2}}\simeq 246\,\text{GeV}$. The Higgs sector is then characterized by eleven independent physical parameters:
@@
In the systematic investigation of Higgs sector phenomenology, it is convenient to adopt specialized parametrizations as follows:
@@
In the basis ($H_{\rm NSM}$, $H_{\rm SM}$, ${\rm Re[S]}$), the mass matrix of CP-even Higgs fields can be written as [28]
@@
Similarly, in the basis ($A_{\rm NSM}$, ${\rm Im[S]}$), the mass matrix of CP-odd Higgs fields can be written as [28]
@@
These correspond to three CP-even states $h_{i}$ (i=1,2,3) and two CP-odd states $a_{j}$ (j=1,2), ordered by increasing mass $m_{h_{1}}<m_{h_{2}}<m_{h_{3}}$ and $m_{a_{1}}<m_{a_{2}}$. In terms of their dominant components, the physical states are also labeled as:
  • $h$: SM-like scalar (125 GeV Higgs)
  • $H$, $A_{H}$: the CP-even and CP-odd doublet-dominated states
  • $h_{s}$, $A_{s}$: the CP-even and CP-odd singlet-dominated states
Critically, the identity of the 125 GeV Higgs depends on the mass hierarchy:
@@
Additionally, the model predicts a pair of charged Higgs bosons:
@@
with mass given by [8, 28]
@@
To connect theoretical inputs with experimental observables, we replace the Lagrangian parameters $\mu,\mu^{\prime},m^{2}_{3},m^{\prime 2}_{S},$ and $\xi^{\prime}$ with physical mass parameters:
  • $m_{A}\equiv\sqrt{M^{2}_{P,11}}$: the mass of heavy MSSM-like CP-odd Higgs boson,
  • $m_{B}\equiv\sqrt{M^{2}_{S,33}}$: the mass of CP-even singlet Higgs boson,
  • $m_{C}\equiv\sqrt{M^{2}_{P,22}}$: the mass of CP-odd singlet Higgs boson,
@@
The original Lagrangian parameters then become:
@@
Critically, the neutral Higgs sector is fully specified by eight parameters: $\tan\beta$, $\lambda$, $A_{\lambda}$, $m_{A}$, $m_{B}$, $m_{C}$, $m_{N}$, and $\mu_{tot}$. The remaining parameters $\kappa$, $A_{\kappa}$, and $v_{s}$ exclusively govern triple Higgs couplings [28].
In the light $h_{s}$ scenario, the scalar mass matrix element ${\cal M}^{2}_{S,23}$ is conventionally parameterized as:
@@
where the dimensionless parameter $\delta\equiv[2\mu_{\rm tot}-(A_{\lambda}+m_{N})\sin 2\beta]/(2\mu_{\rm tot})$ quantifies the cancellation between the effective $\mu$-term and SUSY-breaking contributions. This parametrization offers significant advantages: it naturally accommodates larger $\lambda$ couplings while maintaining compatibility with LHC Higgs data through modest $\delta$ values. Consequently, we adopt $\delta$ as an input parameter instead of $A_{\lambda}$ in this work.

### v3

# Constraining the light Higgs bosons in the GNMSSM with recent Higgs data

@@ Abstract
The search for light scalar and pseudoscalar particles not only provides a promising avenue for probing physics beyond the Standard Model, but also enables a strong first-order electroweak phase transition, thereby offering a viable mechanism to explain the observed baryon asymmetry of the universe. In this study, we investigate exotic decay channels of the 125 GeV Standard-Model-like Higgs boson into pairs of light CP-odd ($a_{s}$) or CP-even ($h_{s}$) Higgs bosons within the framework of the Generalized Next-to-Minimal Supersymmetric Standard Model (GNMSSM). A comprehensive parameter space scan is performed using the MultiNest algorithm, incorporating constraints from HiggsSignals-2.6.2, HiggsBounds-5.10.2, and ATLAS experimental searches, under two distinct scenarios(one in which the observed Higgs boson corresponds to the lighter CP-even state ($h_{1}=h_{\text{SM}}$), and another where it is identified with the heavier state ($h_{2}=h_{\text{SM}}$) ). Our results demonstrate that HiggsBounds imposes the most stringent exclusion limits due to its sensitivity to direct searches for non-Standard Model Higgs bosons. In contrast, HiggsSignals can also exclude regions with suppressed exotic branching ratios—such as $Br(h\to a_{s}a_{s}\to\tau\tau bb)\leq 2.5\%$ in the $h_{2}$-SM scenario—by virtue of its ability to probe indirect deviations arising from kinematically enhanced $h_{2}\to h_{s}h_{s}$ decays. Under combined constraints from HiggsTools, the observed Higgs boson must retain at least 93% SM-like component ($V_{h}^{\text{SM}}\geq 0.93$) with no more than 32% singlet admixture ($V_{h}^{\text{S}}\leq 0.32$); in the $h_{2}$ case, the lighter scalar $h_{s}$ exhibits high singlet purity ($V_{h_{s}}^{\text{S}}\geq 0.94$). Furthermore, dark matter (DM) phenomenology indicates that Singlino- or Higgsino-dominated DM is viable in the $h_{1}$ scenario, with dominant annihilation channels including $\tilde{\chi}_{1}^{0}\tilde{\chi}_{1}^{0}\to h_{s}A_{s}$ for Singlino-like DM and chargino co-annihilation for Higgsino-like DM, whereas the $h_{2}$ scenario favors Higgsino-dominated DM.
@@ 1 Introduction
The discovery of the 125-GeV Higgs boson at the Large Hadron Collider (LHC) [1, 2] marked a milestone in particle physics because it confirmed the mechanism of electroweak symmetry breaking (EWSB) that the Standard Model (SM) describes. However, the SM leaves several fundamental problems unsolved, such as the neutrino masses, hierarchy problem, and dark matter (DM) problem. These unresolved issues have motivated exploration of physics beyond the Standard Model (BSM), where extended Higgs sectors and new symmetries often predict additional scalar states.
@@
Supersymmetry (SUSY) is a BSM candidate that offers solutions to address the hierarchy problem and enable gauge coupling unification. The Minimum Supersymmetric Standard Model (MSSM) [3, 4, 5, 6], which is the most economical SUSY realization, has been extensively studied. However, it is challenged by DM constraints, the $\mu$ problem, the little hierarchy problem, and other issues. As a natural and minimal extension, the Next-to-Minimal Supersymmetric Standard Model (NMSSM) [7, 8, 9] introduces a gauge-singlet Higgs superfield. Its scalar component acquires a non-zero vacuum expectation value (vev), which generates an effective $\mu$-term. Mixing between singlet and doublet superfields produces a new CP-even Higgs ($h_{s}$) that mixes with MSSM-like states. This mixing can suppress the lightest Higgs mass while enabling the next-to-lightest Higgs to match the observed 125-GeV signal [10, 11, 12, 13, 14, 15, 16]. This scenario alleviates fine-tuning and offers rich phenomenology, including Higgs-to-Higgs decays [17, 18, 19, 20, 21, 22, 23, 24] or exotic decays that involve neutralinos [25].
Over the past decade, the NMSSM with the $Z_{3}$-symmetric ($Z_{3}$-invariant NMSSM) [7, 8, 9] has drawn significant interest as a simple extension of the MSSM. However, LHC searches for supersymmetric particles and tighter limits from dark matter experiments such as XENON-nT [26], PandaX-4T [27, 28], and LUX-ZEPLIN (LZ) [29, 30] have severely constrained its parameter space. Searches for light Higgs bosons, electroweakinos, and extra scalars have ruled out much of the natural parameter region, especially where the 125 GeV Higgs mass, B-physics, and dark matter density are simultaneously satisfied [31, 32]. For large $\lambda$, predicted spin-independent scattering cross sections approach current detection limits [31, 32]; only small $\lambda$ or phase cancellations allow evasion, but small $\lambda$ complicates achieving the correct Higgs mass without fine-tuning [31, 32]. Thus, the $Z_{3}$-invariant NMSSM survives only in a narrow and fine-tuned region, challenging its original goal of solving the naturalness problem.
To address the aforementioned theoretical limitations of the $Z_{3}$-invariant NMSSM regarding the Higgs boson mass, the superparticle mass spectrum, and its connection to dark matter, researchers have proposed various extended frameworks, among which the General NMSSM (GNMSSM) [33, 34, 35, 36, 37, 38] stands out as a representative example. By adding extra singlet couplings or $Z_{3}$-symmetry-breaking terms, the GNMSSM enhances the flexibility in the Higgs mass matrix and enables lighter CP-even and CP-odd states, which are consistent with LHC data. Moreover, within the framework of the GNMSSM, the lightest supersymmetric particle (LSP) is a dark matter (DM) candidate. Its annihilation mainly proceeds through s-channel exchange of light CP-even and CP-odd Higgs bosons. The coupling strength determines both the dark matter relic abundance from the early universe and the elastic scattering cross-section with nucleons in direct detection experiments. This link connects collider physics and dark matter phenomenology, allowing constraints from high-energy and astrophysical observations to complement each other.
Couplings between these light scalar particles and standard model fermions or gauge bosons are dynamically suppressed, leading to very low production rates in diphoton, ZZ, or WW channels. As a result, these particles are difficult to detect with conventional searches. However, they can still appear indirectly through hidden-sector effects, especially in exotic Higgs decays. In certain regions, the 125 GeV Higgs boson h decays into pairs of light scalars (e.g., $h\to h_{s}h_{s}$ or $a_{s}a_{s}$ ), each of which then decays into b-jets, $\tau$-lepton pairs, or other standard model states. Although these cascade decays have small branching ratios, their low backgrounds make them sensitive probes for new physics. Therefore, ATLAS and CMS have systematically searched for such signatures across multiple LHC runs, targeting final states such as $4b$ [39], $\mu\mu bb$ [40, 41, 42], $\tau\tau bb$ [43, 44], $\gamma\gamma\tau\tau$ [45], and $4\mu$ [46, 47, 48, 49, 50].
This study focuses on the $\tau\tau bb$ final state [44]. Compared to the $4b$ [39], this signature offers enhanced background suppression by exploiting the distinctive decay topology of the $\tau$ lepton, while retaining high signal efficiency. We use the HiggsTools package[51] to constrain the model parameter space, identifying regions consistent with direct search limits and Higgs property measurements. Based on these constraints, we study $h\to h_{s}h_{s}/a_{s}a_{s}$ exotic decays in the GNMSSM framework and evaluate their observability. We have also conducted a preliminary analysis of the properties of dark matter within the allowed parameter region.
@@
This paper is organized as follows. Section 2 introduces the GNMSSM framework and describes in detail its Higgs sector and the key trilinear Higgs couplings. Section 3 outlines our scanning strategy and the corresponding numerical results. Section 4 summarizes our conclusions.
@@ 2.1 Basics of the GNMSSM
The GNMSSM extends the MSSM by introducing a gauge-singlet superfield $\hat{S}$, which carries neither baryon nor lepton number. Thus, the Higgs sector comprises the standard $SU(2)_{L}$ doublets $\hat{H}_{u}=(\hat{H}^{+}_{u},\hat{H}^{0}_{u})$ and $\hat{H}_{d}=(\hat{H}^{0}_{d},\hat{H}^{-}_{d})$ and the singlet $\hat{S}$. Its gauge-invariant superpotential is as follows  [8]:
@@
where $W_{\rm Yukawa}$ contains the MSSM quark and lepton Yukawa interactions. The dimensionless couplings $\lambda$ and $\kappa$ govern the Higgs interactions, which are analogous to the $Z_{3}$-NMSSM. The GNMSSM incorporates explicit $Z_{3}$-symmetry-breaking via parameters $\mu$, $\mu^{\prime}$ and $\xi$. These terms resolve the inherent tadpole problem [52, 8] and cosmological domain-wall problem [53, 54, 55] in the $Z_{3}$-invariant NMSSM. Parameter $\xi$ can be eliminated by singlet field redefinition  [56] and enables us to set $\xi=0$ without loss of generality. $\mu$ and $\mu^{\prime}$ arise from the spontaneous breaking of discrete $R$-symmetries ($\mathbb{Z}^{R}_{4}$ or $\mathbb{Z}^{R}_{8}$) at high scales [53, 57, 58, 56, 59]. These explicit $Z_{3}$-breaking terms modify the neutral Higgs mass spectrum and generate markedly richer phenomenology than the $Z_{3}$-NMSSM and MSSM.
2.2 Higgs sector of the GNMSSM
@@
where $H_{u},H_{d},$ and $S$ are scalar components of the Higgs superfields, and $m^{2}_{H_{u}},m^{2}_{H_{d}},$ and $m^{2}_{S}$ are their supersymmetry-breaking masses. After the electroweak symmetry breaking, the neutral components of the Higgs fields develop non-vanishing vevs:
@@
with $v=\sqrt{v_{u}^{2}+v_{d}^{2}}\simeq 246\,\text{GeV}$. Then, the Higgs sector is characterized by 11 independent physical parameters:
@@
In the investigation of the Higgs sector phenomenology, it is useful to adopt specialized parametrizations as follows:
@@
In the basis of ($H_{\rm NSM}$, $H_{\rm SM}$, $\rm Re[S]$}), the mass matrix of CP-even Higgs fields can be written as: [35]
@@
Similarly, based on ($A_{\rm NSM}$, $\rm Im[S]$}), the mass matrix of CP-odd Higgs fields can be written as [35]
@@
These mass matrices correspond to three CP-even states $h_{i}$ (i=1,2,3) and two CP-odd states $a_{j}$ (j=1,2), which are ordered by increasing mass: $m_{h_{1}}<m_{h_{2}}<m_{h_{3}}$ and $m_{a_{1}}<m_{a_{2}}$. In terms of dominant components, the physical states are labeled as follows:
  • $h$: SM-like scalar (125-GeV Higgs)
  • $H$, $A_{H}$: CP-even and CP-odd doublet-dominated states
  • $h_{s}$, $A_{s}$: CP-even and CP-odd singlet-dominated states
The identity of the 125-GeV Higgs depends on the mass hierarchy:
@@
In addition, the model predicts a pair of charged Higgs bosons:
@@
with the following mass: [8, 35]
@@
To connect the theoretical inputs with experimental observables, we replaced the Lagrangian parameters $\mu,\mu^{\prime},m^{2}_{3},m^{\prime 2}_{S},$ and $\xi^{\prime}$ with physical mass parameters:
  • $m_{A}\equiv\sqrt{M^{2}_{P,11}}$: mass of a heavy MSSM-like CP-odd Higgs boson,
  • $m_{B}\equiv\sqrt{M^{2}_{S,33}}$: mass of a CP-even singlet Higgs boson,
  • $m_{C}\equiv\sqrt{M^{2}_{P,22}}$: mass of a CP-odd singlet Higgs boson,
@@
Then, the original Lagrangian parameters become:
@@
The neutral Higgs sector is fully specified by eight parameters: $\tan\beta$, $\lambda$, $A_{\lambda}$, $m_{A}$, $m_{B}$, $m_{C}$, $m_{N}$, and $\mu_{tot}$. The remaining parameters ($\kappa$, $A_{\kappa}$, and $v_{s}$) exclusively govern triple Higgs couplings [35].
In the light $h_{s}$ scenario, the scalar mass matrix element ${\cal M}^{2}_{S,23}$ is parameterized as:
@@
where the dimensionless parameter $\delta\equiv[2\mu_{\rm tot}-(A_{\lambda}+m_{N})\sin 2\beta]/(2\mu_{\rm tot})$ quantifies the cancellation between effective $\mu$-term and SUSY-breaking contributions. This parametrization offers advantages: it naturally accommodates larger $\lambda$ couplings while remaining compatible with LHC Higgs data through modest $\delta$ values. Thus, we used $\delta$ as an input parameter instead of $A_{\lambda}$ in this work.
In the scenario with very heavy charged Higgs bosons, approximate expressions can be derived for the singlet-dominated states and their mixing parameters [60]:
$m_{h_{s}}^{2}$ $\simeq$ $m_{B}^{2}-\frac{{\cal M}^{4}_{S,13}}{m_{A}^{2}-m_{B}^{2}},\quad m_{A_{s}}^{2}\simeq m_{C}^{2}-\frac{{\cal M}^{4}_{P,12}}{m_{A}^{2}-m_{C}^{2}},\quad\frac{V_{P,A_{s}}^{\rm NSM}}{V_{P,A_{s}}^{\rm S}}=\frac{{\cal M}^{2}_{P,12}}{m_{A_{s}}^{2}-m_{A}^{2}}\simeq 0,$
$\frac{V_{h}^{\rm S}}{V_{h}^{\rm SM}}$ $\simeq$ $\frac{{\cal M}^{2}_{S,23}}{m_{h}^{2}-m_{B}^{2}},\quad V_{h}^{\rm NSM}\sim 0,\quad V_{h}^{\rm SM}\simeq\sqrt{1-\left(\frac{V_{h}^{\rm S}}{V_{h}^{\rm SM}}\right)^{2}}\sim 1,$
$\frac{V_{h_{s}}^{\rm SM}}{V_{h_{s}}^{\rm S}}$ $\simeq$ $\frac{{\cal M}^{2}_{S,23}}{m_{h_{s}}^{2}-m_{h}^{2}},\quad V_{h_{s}}^{\rm NSM}\sim 0,\quad V_{h_{s}}^{\rm S}\simeq\sqrt{1-\left(\frac{V_{h_{s}}^{\rm SM}}{V_{h_{s}}^{\rm S}}\right)^{2}}\sim 1.$ (2.14)
These expressions indicate that $V_{h}^{S}\simeq-V_{h_{s}}^{\rm SM}$ and that all scale with $\lambda$. In the limit of $\lambda\to 0$, the singlet fields decouple from the doublet Higgs fields, and the masses $m_{B}$, $m_{C}$, and $m_{N}$ can be treated as the physical particle masses with high accuracy. Moreover, the singlet masses are independent and may assume small values, as they are only weakly constrained by current experimental data.
2.3 Neutralino sector of the GNMSSM
[…]
$M_{\tilde{\chi}^{0}}=\left(\begin{array}[]{ccccc}M_{1}&0&-m_{Z}\sin\theta_{W}\cos\beta&m_{Z}\sin\theta_{W}\sin\beta&0\\ &M_{2}&m_{Z}\cos\theta_{W}\cos\beta&-m_{Z}\cos\theta_{W}\sin\beta&0\\ &&0&-\mu_{\text{tot}}&-\frac{1}{\sqrt{2}}\lambda v\sin\beta\\ &&&0&-\frac{1}{\sqrt{2}}\lambda v\cos\beta\\ &&&&m_{\text{N}}\end{array}\right),$ (2.15)
where $M_{1}$ and $M_{2}$ denote the gaugino soft-breaking mass parameters, while $s_{W}\equiv\sin\theta_{W}$ and $c_{W}\equiv\cos\theta_{W}$. By diagonalization the mass matrix using the rotation matrix $N$, the physical neutralino mass eigenstates $\tilde{\chi}^{0}_{i}$ are obtained as follows:
$\tilde{\chi}_{i}^{0}=N_{i1}\psi^{0}_{1}+N_{i2}\psi^{0}_{2}+N_{i3}\psi^{0}_{3}+N_{i4}\psi^{0}_{4}+N_{i5}\psi^{0}_{5},$ (2.16)
where the index i=1,2,,5i=1,2,...,5, arranged in ascending order of mass. The lightest neutralino, $\tilde{\chi}_{1}^{0}$, can serve as the DM candidate. In this state, $N_{13}^{2}+N_{14}^{2}$ and $N_{15}^{2}$ correspond to the Higgsino and Singlino components of the physical state $\tilde{\chi}_{1}^{0}$, respectively. The lightest neutralino $\tilde{\chi}_{1}^{0}$ is classified as Higgsino-dominated DM if $(N_{13}^{2}+N_{14}^{2})>0.5$ and Singlino-dominated DM if $N_{15}^{2}>0.5$. In scenarios where gauginos are significantly heavier and the condition $\mu_{tot}^{2}-m_{N}^{2}>>\lambda^{2}v^{2}$ holds, the mass of the Singlino-dominated DM can be approximately given by [35]
$m_{\tilde{\chi}_{1}^{0}}\simeq m_{N}+\frac{\lambda^{2}v^{2}(m_{\tilde{\chi}_{1}^{0}}-\mu_{tot}\sin 2\beta)}{2(m_{\tilde{\chi}_{1}^{0}}^{2}-\mu_{tot}^{2})}\simeq m_{N}$ (2.17)